OpenStudy (anonymous):

Given the following differential equation find the solution that passes through (1,1). You may use separation of variables. dy/dt = t / (y^2 * sqrt(1+t^2))

OpenStudy (amistre64):

[S] y^2 dy =[S] t/sqrt(1+t^2) dt

OpenStudy (amistre64):

(1/3)y^3 = (1/2)(sqrt(1+t^2)) is that right :)

OpenStudy (amistre64):

i think that ''1/2'' cancels ....

OpenStudy (amistre64):

(1/3)y^3 = sqrt(1+t^2) +c 1,1 seems like coordinantes, but are they t=1 and y=1?

OpenStudy (amistre64):

(1/3) = sqrt(2) +c c = sqrt(2)/3 (1/3) y^3 = sqrt(1+t^2) + sqrt(2)/3....thats the best i got....

OpenStudy (anonymous):

Thank you very much! c = (1/3) - sqrt(2) The final answer is: (3(sqrt(1+t^2)+((1/3)-sqrt(2))))^(1/3) = y(t) Thanks!

OpenStudy (amistre64):

youre welcome :) just glad I was right lol

OpenStudy (amistre64):

er...kinda right